OpenStudy (haleyelizabeth2017):

Graph the function. What is the amplitude, period, phase shift, and vertical shift for y=4 sin (3theta-1/3 pi) +1 I believe that: Amp=4 period=2pi/3 phase shift=1/3pi vertical shift=1

Directrix (directrix):

This chart may help.

OpenStudy (haleyelizabeth2017):

so 1/9pi is the phase shift?

OpenStudy (haleyelizabeth2017):

(1/9)pi **

Directrix (directrix):

Is this: 1/3 pi 1/3 of pi or 1 over 3 pi as in 1/(3 pi)

OpenStudy (haleyelizabeth2017):

it's (1/3)pi

OpenStudy (haleyelizabeth2017):

Sorry about it being unclear...ha...ha

Directrix (directrix):

I wrote the function this way for comparison with the chart: y = 4 sin ( 3Θ - π/3 ) + 1 I agree with you on these three: Amp=4 period=2pi/3 vertical shift=1

Directrix (directrix):

Now, the phase shift.

OpenStudy (haleyelizabeth2017):

c=(1/3)pi b=3 (1/3)pi/3 = (1/3)pi * (1/3) (1/9)pi

Directrix (directrix):

-C/B where C = - π/3 and B = 3 - ( - π/3 ) / 3 = π/9

Directrix (directrix):

phase shift = π/9 which is what you got.

OpenStudy (haleyelizabeth2017):

:)